# Families of cyclic curve coverings with maximal monodromy

### v1

@@ 1. Introduction
where $m_{\chi}$ is the multiplicity of the $\chi$-eigenspace for the action of $G$ on $H^{1,0}(C_{b})$ (and similarly for the conjugate $\bar{\chi}$). Taking the adjoint in (1.1), one forgets the factors coming from 0-dimensional eigenspaces. In the following, we consider families with no repeating factors in the monodromy, meaning that, for any two characters $\chi_{1}$ and $\chi_{2}$ with $m_{\chi_{i}},m_{\bar{\chi}_{i}}\neq 0$, the multiplicities satisfy $\{m_{\chi_{1}},\,m_{\bar{\chi}_{1}}\}\neq\{m_{\chi_{2}},\,m_{\bar{\chi}_{2}}\}$. In other words, a family has no repeating factors in the monodromy if all the factors appearing in $(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad}$ are mutually distinct. Our first result is the following.
@@
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings with no repeating factors in the monodromy. Then the monodromy group is maximal, namely
$\operatorname{Mon}^{0,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad}.$
In particular, if $g^{\prime}=g(C^{\prime})\geq 2$, we have
$\operatorname{Mon}^{0}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der}.$
(See Theorem 4.7). Maximality results for the monodromy of families of algebraic varieties are widely studied in algebraic geometry. Regarding $G$-coverings of curves, we mention e.g. [4, 3, 19, 17]. In particular we point out that, for $g^{\prime}\geq 4$, Theorem A partially recovers [17, Theorem 1.3], where the authors prove a strong maximality result for the monodromy of families of any Galois covers, requiring some lower bound on the genus of the base curve $C^{\prime}$. By studying the adjoint monodromy group $\operatorname{Mon}^{0,ad}$, we ignore the contribution of 0-dimensional eigenspaces and, thus, we can take into account lower genus $g^{\prime}$, too. In fact, our proof proceeds by induction on the genus $g^{\prime}$. The case $g^{\prime}=0$ follows from [26]. The induction step is based on a degeneration argument and requires the study of Prym varieties associated with certain admissible coverings, which is carried out in Theorem 3.3. As a consequence, under the assumption of cyclic coverings with no repeating factors, our argument applies to all genera $g^{\prime}$, and in the case $g^{\prime}\geq 4$ it provides an alternative proof of the maximality of the monodromy. Notice that, if $g^{\prime}\geq 2$, no 0-dimensional eigenspace appears and, hence, our result concerns the full monodromy group $\operatorname{Mon}^{0}$.
@@
The maximality of the monodromy is a powerful property and it has strong consequences on the geometry of the moduli image of our families. In particular, our result has interesting applications to the Coleman-Oort conjecture. Denote by $\mathcal{M}_{g}$ the moduli space of smooth projective curves of genus $g$ over $\mathbb{C}$, by $\mathcal{A}_{g}$ the moduli space of $g$-dimensional principally polarized abelian varieties, and let $j:\mathcal{M}_{g}\to\mathcal{A}_{g}$ be Torelli map sending curves to their Jacobians. The Coleman-Oort conjecture concerns the interaction between the image of the Torelli map and the special subvarieties of $\mathcal{A}_{g}$, which are defined as Hodge loci associated with the natural rational variation of Hodge structure on $\mathcal{A}_{g}$. Special subvarieties are totally geodesic, namely they are linear with respect to the locally symmetric metric induced from the Siegel space. The conjecture predicts that for $g\gg 0$, there are no positive-dimensional special subvarieties $S\subset\mathcal{A}_{g}$ such that $S$ is generically contained in the Torelli locus, i.e. $S\subset\overline{j(\mathcal{M}_{g})}$ and $S\cap j(\mathcal{M}_{g})\neq\emptyset$. However, for low genus, counterexamples are known: there exist positive dimensional special subvarieties $S\subset\mathcal{A}_{g}$ that lie generically in the Torelli locus. All known examples occur for $g\leq 7$, and arise from families of Galois covers of curves. Several such examples have been constructed and studied in the literature; see for instance [8, 24, 22, 23, 10, 11, 13, 16, 20, 25, 27]; we recall their main construction in section 5. The classification of such examples in case of cyclic covering of $\mathbb{P}^{1}$ is due to Moonen ([22]). The complete list of all currently known examples of special families of Galois coverings is provided in [10, 11]. Our second result provides new evidence for the Coleman-Oort conjecture.
@@
(See Theorem 5.5). More precisely, we prove that the only positive dimensional totally geodesic families of cyclic coverings with no repeating factors in the monodromy are among the known ones (namely, among the ones appearing in [10, 11]). Moreover, we observe that, even though assuming no repeating factors appears as a strong condition on the monodromy, we have:
@@
(See Theorem 5.8). Theorem B follows combining Theorem A with the fact that the smallest totally geodesic subvariety containing the moduli image of a family is given by (the image in $\mathcal{A}_{g}$) of the orbit of $\operatorname{Mon}^{0,ad}_{\mathbb{R}}$. When $g^{\prime}=0$, our results recover Moonen’s classification of families of cyclic covers of the projective line under the assumption that the monodromy has no repeating factors [22]. Furthermore, this assumption allows us to treat the case $g^{\prime}>0$ as well. We point out that, along with the proof, we also show that families with no repeating factors in the monodromy are special if and only if they are totally geodesic (see Proposition 5.4).
@@ 2. Preliminaries on families of $G$-curves
(2.1) $m_{\chi_{i}}=(g^{\prime}-1)+\sum_{i=1}^{r}\frac{[i\theta_{i}]_{d}}{d},$
where $\chi_{i}$ is the representation associated with the eigenvalue $\xi^{i}$, being $\xi$ a primitive $d$-th root of the unity, and $a]_{d}$ denotes the unique representative of $a\in\mathbb{Z}/d\mathbb{Z}$ in $\{0,...,d-1\}$.
@@ 3. An admissible covering and its Prym variety
More concretely, let $p_{1},p_{2}\in D^{\nu}$ be the two preimages of the node of $D$. These points are the branch points of the normalized cover $f_{D}^{\nu}$, with local monodromies $1$ and 1-1, respectively. The admissible cover $f_{D}$ is obtained from its normalization $f_{D}^{\nu}$ by gluing the points $p_{1}$ and $p_{2}$ in $D^{\nu}$ to form the node of $D$, and simultaneously gluing their corresponding preimages in $\tilde{D}\nu$. Note that we have $g(\tilde{D}^{\nu})=g-1$ and $g(D^{\nu})=g^{\prime}-1$.
We now study the Prym variety associated with the covering $f_{D}$, inspired by [2, 9, 18].
At the node $s$ of $\tilde{D}$, we make the usual identification:
@@ 4. Maximal monodromy
Let $\operatorname{Mon}^{0}\subset\operatorname{GL}(H^{1}(C_{b},{\mathbb{Q}}))$ denote the connected monodromy group, and $\operatorname{Hg}\subset\operatorname{GL}(H^{1}(C_{b},{\mathbb{Q}}))$ the generic Hodge group of the VHS. By a theorem of André [1, Theorem 1], $\operatorname{Mon}^{0}$ is a semisimple normal subgroup of $\operatorname{Hg}$. More precisely, we have
@@
(4.3) $\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G}\simeq\prod_{(\chi,\bar{\chi})\in X(G,\mathbb{R})}\operatorname{Sp}(H^{1}(C_{b},\mathbb{R})_{\chi})^{G}\simeq\prod_{(\chi,\bar{\chi}),\chi\neq\bar{\chi}}\operatorname{U}(m_{\chi},m_{\bar{\chi}})\times\prod_{(\chi,\bar{\chi}),\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
We denote by $\operatorname{Mon}^{0}(\chi)$ the projection of $\operatorname{Mon}^{0}_{\mathbb{R}}$ onto the $\chi$-factor of the decomposition. If $\chi\neq\bar{\chi}$, we have $\operatorname{Mon}^{0}(\chi)\subseteq\operatorname{SU}(m_{\chi},m_{\bar{\chi}})$, whether for $\chi=\bar{\chi}$, then $\operatorname{Mon}^{0}(\chi)\subseteq\operatorname{Sp}_{2m_{\chi}}.$ Thus, we have
(4.4) $\operatorname{Mon}^{0}_{\mathbb{R}}\hookrightarrow\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
@@
In other words, a family has no repeating factors in the monodromy if all the factors appearing in $(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad}$ are mutually distinct. Our main theorem uses an induction argument on the genus $g^{\prime}$ of the base curve $C^{\prime}_{b}$. The base of the induction, i.e. the case with $g^{\prime}=0$, is the following result:
@@
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings of the line with no repeating factors in the monodromy. Then the monodromy is maximal, that is
$\operatorname{Mon}^{0,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad}.$
This result is proved in [26, Theorem 1.7]. A fundamental ingredient, due to Rohde [24], is that for cyclic coverings of the line, whenever the multiplicities $m_{\chi}$ and $m_{\bar{\chi}}$ are nonzero, one has $\operatorname{Mon}^{0}(\chi)=\operatorname{SU}(m_{\chi},m_{\bar{\chi}})$ if $\chi\neq\bar{\chi}$, and $\operatorname{Mon}^{0}(\chi)=\operatorname{Sp}_{2m_{\chi}}$ if $\chi=\bar{\chi}$.
@@
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings with no repeating factors in the monodromy. Then the monodromy is maximal, that is
(4.5) $\operatorname{Mon}^{0,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad},$
and, in particular, if $g^{\prime}\geq 2$
(4.6) $\operatorname{Mon}^{0}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der}.$
@@
where the first term in the product comes from the trivial representation. As $C_{b}^{\prime}$ moves in the whole moduli space $\mathcal{M}_{g^{\prime}}$, we have $\operatorname{Mon}^{0}(JC^{\prime}_{b})=\operatorname{Sp}_{2g^{\prime}}$. Hence, to prove maximality it is enough to focus on $\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$. We first observe that we have
$\operatorname{Mon}^{0,ad}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))\subseteq\prod_{\chi\neq\bar{\chi},\\ \ m_{\chi}m_{\bar{\chi}}\neq 0}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi},\\ m_{\chi}\neq 0}\operatorname{Sp}(2m_{\chi},\mathbb{R}),$
and using a degeneration, we can show that we actually have equality. Indeed, as in 4.2, we degenerate to the admissible cover $\tilde{D}\rightarrow D$ whose normalization $\tilde{D}^{\nu}\rightarrow D^{\nu}$ is endowed with monodromy $\Theta^{\nu}=(1,-1,\theta_{1},\dots,\theta_{s})$. By Theorem 3.3, we have $\operatorname{Prym}(\tilde{D},D)\sim\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu})$. By Lemma 4.6, the covering $\tilde{D}^{\nu}\rightarrow D^{\nu}$ has
@@
for all $\chi_{i},\chi_{j}\in X(G)$ different from the trivial representation, $i\neq j$ with $m^{\nu}_{\chi_{i}}m^{\nu}_{\bar{\chi}_{i}}\neq 0$. In other words, $\tilde{D}^{\nu}\rightarrow D^{\nu}$ also has no repeating factors in the monodromy. Since $g(D^{\nu})=g^{\prime}-1$, the induction hypothesis implies that $\operatorname{Mon}^{0}(\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu}))$ is maximal. Moreover we have $\operatorname{Mon}^{0,ad}(\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu}))=\operatorname{Mon}^{0,ad}(\operatorname{Prym}(\tilde{D},D))\subseteq\operatorname{Mon}^{0,ad}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$, which then gives maximality for $\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$, i.e. (4.5). Using Chevalley-Weil formula 2.1, we get that no 0-dimensional eigenspace appears if $g^{\prime}\geq 2$. Thus, (4.6) follows. ∎

### v2

@@ 1. Introduction
where $m_{\chi}$ is the multiplicity of the $\chi$-eigenspace for the action of $G$ on $H^{1,0}(C_{b})$ (and similarly for the conjugate $\bar{\chi}$). In the following, we consider families with no repeating factors in the monodromy, meaning that, for any two characters $\chi_{1}$ and $\chi_{2}$ with $m_{\chi_{i}},m_{\bar{\chi}_{i}}\neq 0$, the multiplicities satisfy $\{m_{\chi_{1}},\,m_{\bar{\chi}_{1}}\}\neq\{m_{\chi_{2}},\,m_{\bar{\chi}_{2}}\}$. Our first result is the following.
@@
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings with no repeating factors in the monodromy with $g^{\prime}=g(C_{b}^{\prime})\geq 1$. Then the monodromy group is maximal, namely
$\operatorname{Mon}^{0}_{\mathbb{R}}=(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{der}.$
(See Theorem 4.7). Maximality results for the monodromy of families of algebraic varieties are widely studied in algebraic geometry. Regarding $G$-coverings of curves, we mention e.g. [4, 3, 19, 17]. In particular we point out that, for $g^{\prime}\geq 4$, Theorem A partially recovers [17, Theorem 1.3], where the authors prove a strong maximality result for the monodromy of families of any Galois covers, requiring some lower bound on the genus of the base curve $C^{\prime}$. Our argument allows to take into account lower genus $g^{\prime}$, too. In fact, our proof proceeds by induction on the genus $g^{\prime}$. In the case $g^{\prime}=0$, we show that $(\operatorname{Sp}^{G})^{pos}\subseteq\operatorname{Mon}^{0}_{\mathbb{R}}$, where $(\operatorname{Sp}^{G})^{pos}$ involves only the contributions in $(\operatorname{Sp}^{G})^{der}$ coming from positive-dimensional eigenspaces (see Theorem 4.4). The induction step is based on a degeneration argument and requires the study of Prym varieties associated with certain admissible coverings, which is carried out in Theorem 3.3. If $g^{\prime}\geq 1$, we show that $(\operatorname{Sp}^{G})^{der}=(\operatorname{Sp}^{G})^{pos}$, yielding Theorem A. As a consequence, under the assumption of cyclic coverings with no repeating factors, our argument applies to all genera $g^{\prime}\geq 1$, and in the case $g^{\prime}\geq 4$ it provides an alternative proof of the maximality of the monodromy discussed in [17].
@@
The maximality of the monodromy is a powerful property and it has strong consequences on the geometry of the moduli image of our families. In particular, our result has interesting applications to the Coleman-Oort conjecture. Denote by $\mathcal{M}_{g}$ the moduli space of smooth projective curves of genus $g$ over $\mathbb{C}$, by $\mathcal{A}_{g}$ the moduli space of $g$-dimensional principally polarized abelian varieties, and let $j:\mathcal{M}_{g}\to\mathcal{A}_{g}$ be Torelli map sending curves to their Jacobians. The Coleman-Oort conjecture concerns the interaction between the image of the Torelli map and the special subvarieties of $\mathcal{A}_{g}$, which are defined as Hodge loci associated with the natural rational variation of Hodge structure on $\mathcal{A}_{g}$. Special subvarieties are totally geodesic, namely they are linear with respect to the locally symmetric metric induced from the Siegel space. The conjecture predicts that for $g\gg 0$, there are no positive-dimensional special subvarieties $S\subset\mathcal{A}_{g}$ such that $S$ is generically contained in the Torelli locus, i.e. $S\subset\overline{j(\mathcal{M}_{g})}$ and $S\cap j(\mathcal{M}_{g})\neq\emptyset$. However, for low genus, counterexamples are known: there exist positive dimensional special subvarieties $S\subset\mathcal{A}_{g}$ that lie generically in the Torelli locus. All known examples occur for $g\leq 7$, and arise from families of Galois covers of curves. Several such examples have been constructed and studied in the literature; see for instance [8, 23, 22, 10, 11, 13, 16, 20, 24, 26]; we recall their main construction in section 5. The classification of such examples in case of cyclic covering of $\mathbb{P}^{1}$ is due to Moonen ([22]). The complete list of all currently known examples of special families of Galois coverings is provided in [10, 11]. Our second result provides new evidence for the Coleman-Oort conjecture.
@@
(See Theorem 5.5). More precisely, we prove that the only positive dimensional totally geodesic families of cyclic coverings with no repeating factors in the monodromy are among the known ones (namely, among the ones appearing in [10, 11]). In other words, we provide a classification of all totally geodesic families of cyclic coverings with no repeating factors in the monodromy for all $g^{\prime}\geq 0$. Theorem B follows combining Theorem A with the fact that the smallest totally geodesic subvariety containing the moduli image of a family is given by (the image in $\mathcal{A}_{g}$) of the orbit of $\operatorname{Mon}^{0,ad}_{\mathbb{R}}$. We point out that, along with the proof, we also show that families with no repeating factors in the monodromy are special if and only if they are totally geodesic (see Proposition 5.4). Moreover, we observe that, even though assuming no repeating factors appears as a strong condition on the monodromy, we have:
@@
(See Theorem 5.8).
@@ 2. Preliminaries on families of $G$-curves
(2.1) $m_{\chi_{i}}=(g^{\prime}-1)+\sum_{i=1}^{r}\frac{[i\theta_{i}]_{d}}{d}+\varepsilon,$
where $\chi_{i}$ is the representation associated with the eigenvalue $\xi^{i}$, $\xi$ is a primitive $d$-th root of the unity, $a]_{d}$ denotes the unique representative of $a\in\mathbb{Z}/d\mathbb{Z}$ in $\{0,...,d-1\}$, and $\varepsilon$ is $1$ when $\chi$ is the trivial representation and 0 otherwise.
@@ 3. An admissible covering and its Prym variety
More concretely, let $p_{1},p_{2}\in D^{\nu}$ be the two preimages of the node of $D$. These points are the branch points of the normalized cover $f_{D}^{\nu}$, with local monodromies $1$ and 1-1, respectively. The admissible cover $f_{D}$ is obtained from its normalization $f_{D}^{\nu}$ by gluing the points $p_{1}$ and $p_{2}$ in $D^{\nu}$ to form the node of $D$, and simultaneously gluing their corresponding preimages in $\tilde{D}\nu$. Note that we have $g(\tilde{D}^{\nu})=g-1$ and $g(D^{\nu})=g^{\prime}-1$. An example of such an admissible cover is presented in Figure 1.
d:1p1p2
Figure 1.
Example of a $4:1$ admissible cover (left), where the sheets are glued according to (3.1), and its normalization (right).
We now study the Prym variety associated with the covering $f_{D}$, inspired by [2, 9, 18]. At the node $s$ of $\tilde{D}$, we make the usual identification:
@@ 4. Maximal monodromy
Let $\operatorname{Mon}^{0}\subset\operatorname{GL}(H^{1}(C_{b},{\mathbb{Q}}))$ denote the connected monodromy group, and $\operatorname{Hg}\subset\operatorname{GL}(H^{1}(C_{b},{\mathbb{Q}}))$ the generic Hodge group of the VHS. By a theorem of André [1, Theorem 1], $\operatorname{Mon}^{0}$ is a semisimple normal subgroup of $\operatorname{Hg}^{der}$. More precisely, we have
@@
(4.3) $\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G}$ $\simeq\prod_{(\chi,\bar{\chi})\in X(G,\mathbb{R})}\operatorname{Sp}(H^{1}(C_{b},\mathbb{R})_{\chi})^{G}$
$\simeq\prod_{(\chi,\bar{\chi}),\chi\neq\bar{\chi}}\operatorname{U}(m_{\chi},m_{\bar{\chi}})\times\prod_{(\chi,\bar{\chi}),\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
Thus, we have
(4.4) $\operatorname{Mon}^{0}_{\mathbb{R}}\hookrightarrow(\operatorname{Sp}^{G})^{der}=\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
We denote by $\operatorname{Mon}^{0}(\chi)$ the projection of $\operatorname{Mon}^{0}_{\mathbb{R}}$ onto the $\chi$-factor of the decomposition. If $\chi\neq\bar{\chi}$, we have $\operatorname{Mon}^{0}(\chi)\subseteq\operatorname{SU}(m_{\chi},m_{\bar{\chi}})$, whether for $\chi=\bar{\chi}$, then $\operatorname{Mon}^{0}(\chi)\subseteq\operatorname{Sp}_{2m_{\chi}}.$
@@
From now on we set
(4.5) $(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{pos}:=\prod_{\chi\neq\bar{\chi},\,m_{\chi}m_{\bar{\chi}}\neq 0}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi},\,m_{\chi}\neq 0}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
Hence, a family has no repeating factors in the monodromy if in $(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{pos}$ all the factors are mutually distinct. Our main theorem uses an induction argument on the genus $g^{\prime}$ of the base curve $C^{\prime}_{b}$. The base of the induction, i.e. the case with $g^{\prime}=0$, is the following result:
@@
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings of the line with no repeating factors in the monodromy. Then the monodromy satisfies
$(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{pos}\subseteq\operatorname{Mon}^{0}_{\mathbb{R}}.$
Proof.
This result is proved in [25, Theorem 1.7], for the reader convenience we sketch the proof. The first ingredient, due to Rohde [23], is that for cyclic coverings of the line, whenever the multiplicities $m_{\chi}$ and $m_{\bar{\chi}}$ are nonzero, one has $\operatorname{Mon}^{0}(\chi)=\operatorname{SU}(m_{\chi},m_{\bar{\chi}})$ if $\chi\neq\bar{\chi}$, and $\operatorname{Mon}^{0}(\chi)=\operatorname{Sp}_{2m_{\chi}}$ if $\chi=\bar{\chi}$.
By definition $\operatorname{Mon}^{0}_{\mathbb{R}}$ is a semisimple group contained in $(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{der}$, and by what recalled above, it projects surjectively onto each factor of $(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{pos}$. Thus, since we are assuming that no repeating factor in $(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{pos}$ appears, no diagonal embedding is allowed, and thus we get the claimed inclusion.
@@
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings with no repeating factors in the monodromy. Then the monodromy satisfies
(4.6) $(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{pos}\subseteq\operatorname{Mon}^{0}_{\mathbb{R}}.$
Moreover, if $g^{\prime}\geq 1$, the monodromy is maximal, namely
(4.7) $\operatorname{Mon}^{0}_{\mathbb{R}}=(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{der}.$
@@
where the first term in the product comes from the trivial representation. As $C_{b}^{\prime}$ moves in the whole moduli space $\mathcal{M}_{g^{\prime}}$, we have $\operatorname{Mon}^{0}(JC^{\prime}_{b})=\operatorname{Sp}_{2g^{\prime}}$. Hence, it is enough to focus on $\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$. Using a degeneration, we can show that
(4.8) $\prod_{\chi\neq\bar{\chi},m_{\chi}m_{\bar{\chi}}\neq 0}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi},m_{\chi}\neq 0}\operatorname{Sp}(2m_{\chi},\mathbb{R})\subseteq\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime})).$
Indeed, as in 4.2, we degenerate to the admissible cover $\tilde{D}\rightarrow D$ whose normalization $\tilde{D}^{\nu}\rightarrow D^{\nu}$ is endowed with monodromy $\Theta^{\nu}=(1,-1,\theta_{1},\dots,\theta_{s})$. By Theorem 3.3, we have $\operatorname{Prym}(\tilde{D},D)\sim\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu})$. By Lemma 4.6, the covering $\tilde{D}^{\nu}\rightarrow D^{\nu}$ has
@@
for all $\chi_{i},\chi_{j}\in X(G)$ different from the trivial representation, $i\neq j$ with $m^{\nu}_{\chi_{i}}m^{\nu}_{\bar{\chi}_{i}}\neq 0$. In other words, $\tilde{D}^{\nu}\rightarrow D^{\nu}$ also has no repeating factors in the monodromy. Since $g(D^{\nu})=g^{\prime}-1$, the induction hypothesis implies that
$\prod_{\chi\neq\bar{\chi},\\ \ m_{\chi}m_{\bar{\chi}}\neq 0}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi},\\ m_{\chi}\neq 0}\operatorname{Sp}(2m_{\chi},\mathbb{R})\subseteq\operatorname{Mon}^{0}(\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu})).$
Moreover we have $\operatorname{Mon}^{0}(\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu}))=\operatorname{Mon}^{0}(\operatorname{Prym}(\tilde{D},D))\subseteq\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$, which then implies (4.8), i.e. (4.6) holds.
To complete the proof, we observe that, using Chevalley-Weil formula 2.1, no 0-dimensional eigenspace appears if $g^{\prime}\geq 2$. Thus,
(4.9) $(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{der}=(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{pos}\subseteq\operatorname{Mon}^{0}_{\mathbb{R}}\subseteq(\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G})^{der},$
yielding the equality (4.7) for all $g^{\prime}\geq 2$. On the other hand, when $g^{\prime}=1$, again Chevalley-Weil formula implies that, if there exists a representation $\chi_{i}$ with $m_{\chi_{i}}=0$, then $m_{{\bar{\chi}}_{i}}=0$ as well. This implies that the $\chi_{i}$-factor in $\operatorname{Sp}^{G}$ is trivial. Hence, the equality (4.9) holds when $g^{\prime}=1$ too, and thus we conclude. ∎
Remark 4.8.
It can be shown that the maximality of the monodromy implies the Putman-Wieland conjecture. This concerns families of Galois coverings $\mathcal{C}\rightarrow B$ where $B\rightarrow\mathcal{M}_{g^{\prime},r}$ is dominant, and predicts that $\text{Jac}_{\mathcal{C/B}}$ does not have an isotrivial isogeny factor (i.e. no fixed part). As a consequence, we observe that the bound $g^{\prime}\geq 1$ in Theorem 4.7 for the equality in (4.7) is sharp. In particular, when $g^{\prime}=0$ the inclusion (4.6) is the best we can hope for. This comes from the fact that, when $g^{\prime}=0$ there do exist examples of families with a fixed part in the Jacobians, hence, $\operatorname{Mon}^{0}_{\mathbb{R}}\neq(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der}$. For instance, the Jacobians of the family
$C_{b}:y^{4}=x(x-1)(x-t),\;t\in\mathbb{P}^{1}\smallsetminus\{0,1,\infty\},$
up to isogeny, have a constant elliptic factor.